Solve the following system of nonlinear equations. Choose all answers that apply. x^2+y^2=3 x^2+y=0
Answer by ikleyn(52873)

.
Solve the following system of nonlinear equations. Choose all answers that apply. x^2+y^2=3 x^2+y=0
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

From the second equation express x^2 = y and substitute "-y" into the first equation instead of x^2. 

You will get single quadratic equation for y:

-y + y^2 = 3,   or

y^2 - y - 3 = 0.

Can you complete the assignment from this point on your own?

There are other lessons on systems of equations that are not linear:
    - Solving the system of algebraic equations of degree 2 and degree 1
    - Solving the system of algebraic equations of degree 2
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of equations that are not linear".